Opened 9 years ago

Closed 3 years ago

#350 closed flight processing (fixed)

EUFAR10/02 (AIRES-CZM), 199b/2010, Santander

Reported by: jb362 Owned by: knpa
Priority: alpha 4 high Milestone: 2010 data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by anch)

Data location: ~arsf/arsf_data/2010/flight_data/western_europe/EUFAR10_02-2010_199b-Santander

Data arrived from ARSF via network transfer on 26/07/2010.

Scientific objectives:

1 Development of a reliable tool to estimate coastal water quality indicators by using remote sensing techniques, both airborne and satellite.

1.1. Development of a set of semi-empirical algorithms based on hyperspectral airborne and in-situ data for estimating the concentrations of the water constituents;
1.2. Migration of the developed algorithms into the satellite workflow.

  1. Implementation of an automatic warning system based on satellite imagery processing;
  1. Inclusion of the information obtained by means of remote sensing in the dataset used for the calibration of regional hydrodynamic models.

Priority: a4m - awaiting confirmation

PI: Castillo-Lopez

Sensors: ¶

  • Eagle (27/08/2010)
  • Hawk (21/12/2010)
  • Leica LIDAR (21/12/2010)
  • RCD Digital photography (27/08/2010)

Change History (49)

comment:1 Changed 9 years ago by jb362

  • Owner set to jb362
  • Status changed from new to accepted

comment:2 Changed 9 years ago by jb362

  • Owner changed from jb362 to iopa
  • Status changed from accepted to assigned

comment:3 Changed 9 years ago by iopa

  • Owner changed from iopa to jb362

comment:4 Changed 9 years ago by jb362

  • Priority changed from alpha 4 medium to alpha 4 high

comment:5 Changed 9 years ago by anch

  • Owner changed from jb362 to anch
  • Status changed from assigned to accepted

I'll start processing the Applanix for this now.

comment:6 Changed 9 years ago by anch

Applanix and IPAS for this project completed.

comment:7 Changed 9 years ago by anch

Because this project is not in the UK, and making a DEM using STRM data is a pain, I have started LiDAR processing which I will eventually use to produce a DEM to process the hyperspectral with.

comment:8 Changed 9 years ago by jb362

  • Description modified (diff)

comment:9 Changed 9 years ago by anch

  • Description modified (diff)

LiDAR has lots of noise (mainly over ocean) and may take a while to classify, so I have put the hyperspectral back on the processing list in case anyone is looking for a hyperspectral project to get on with.

comment:10 Changed 9 years ago by jb362

Starting hyperspectral processing.

comment:11 Changed 9 years ago by iopa

  • Description modified (diff)

Processing camera.

comment:12 Changed 9 years ago by jb362

  • Summary changed from EUFAR10/01, 199b/2010, Santander to EUFAR10/02, 199b/2010, Santander

comment:13 Changed 9 years ago by iopa

Checking hyperspectral.

comment:14 Changed 9 years ago by iopa

Hyperspectal delivery check report:

1) rename lev1's and screenshots to contain the lettered julian day 199b.

Otherwise all's well.

comment:15 Changed 9 years ago by jb362

Delivery check fixes complete.

comment:16 Changed 9 years ago by jb362

Checking camera

comment:17 Changed 9 years ago by jb362

All good, but again PI needs to be Dr. Castillo-Lopez

comment:18 Changed 9 years ago by iopa

Camera is ready to deliver.

comment:19 Changed 9 years ago by jb362

Camera and Eagle data delivered.

comment:20 Changed 9 years ago by jb362

  • Description modified (diff)

comment:21 Changed 9 years ago by knpa

  • Description modified (diff)

Data moved from overflow to repository.
Can now be r-synced.

comment:22 Changed 9 years ago by benj

Have downloaded 1s RINEX data from University of Cantabria and copied to repository. Relevant section of email from PI was as follows:

SANTANDER and TORRELAVEGA stations are in relation with the calibration site "Santander1" and the LAREDO station for the area of validation (Santander 2). Santander 1 data taken on 18th July and Santander2  data taken on 19th July.
In the corresponding tabs appear links to download RINEX data.
 
LINKS of interest:
 
http://www.gnss.unican.es/estaciones/cant.asp
 
http://www.gnss.unican.es/estaciones/trvg.asp
 
http://www.gnss.unican.es/estaciones/lare.asp
 
ftp://ftpgnss.digteg.unican.es/rinex
 
Access to real time corrections using the standard protocol NTRIP (Networked Transport of RTCM vía Internet Protocol):
 
URL 	http://ntrip.unican.es:2101/	 
IP:    	193.144.202.127	 
Port: 	21 
User: 	anonimous
 
link: http://www.gnss.unican.es/servicios/services.asp

comment:23 Changed 9 years ago by benj

Note I believe the Santander station mentioned with code CANT is the data I have downloaded from REIN. This is because if you go to the first link given the title of the web page is "Estación de Reinosa"

comment:24 Changed 9 years ago by knpa

Appears to be an issue with the bandset labels in the header files; they start at 900nm which is lower than it should be. Possibly incorrect bandset selected or none specified. Will leave this until end of the year and new Hawk calibration.

comment:25 Changed 9 years ago by benj

  • Description modified (diff)

comment:26 Changed 9 years ago by iopa

LiDAR data has been checked;

readme contains the DEM section title twice
in the data quality section it is written elipsoid with one l

dem/README...txt needs correcting for SRTM references

the data seems correct, ready to deliver.

comment:27 Changed 9 years ago by emca

Eagle and Hawk SCT's

Line Eagle Hawk
1 -0.05 0.02
2 0.08 -0.06
3 0.03 0.01
4 0.05 -0.03
5 -0.05 -0.01
6 0.01 0.01
7 0.02 -0.03
8 -0.01 -0.01
9 -0.03 0.02
10 -0.07 0.02
11 0.01 -0.01
12 0.03 0.04
13 0.35 0.52
14 0.08 0.05
15 0.04 0.04
16 0.08 -0.05
17 -0.06 0.00
18 -0.01 0.00
19 0.00 -0.01

I used a LiDAR intensity image and the Eagle data to find the SCT's for the Hawk as the lines were not overlapping and most of them did not contain sufficient wobbles to correct the timing. Hawk line 13 had no sync. Lines 11 & 19 had bends so were processed using -l 0 7500 and -l 750 17968 respectively, cutting off ~ 200 lines of data.
Hawk delivery has been created.

comment:28 Changed 9 years ago by anch

Hawk Delivery checked.

Looks fine to deliver.

comment:29 Changed 9 years ago by emca

Hawk lev1 files reprocessed and saved in delivery folder

comment:30 Changed 9 years ago by anch

Hawk lev1 re-checked: Ready for delivery

comment:31 Changed 9 years ago by anch

  • Description modified (diff)

Hawk and Lidar delivered 21/12/2010

comment:32 Changed 9 years ago by emca

  • Component changed from Processing: general to Archiving

All data has been r-sync'd to
~arsf/arsf_data/2010/flight_data/western_europe/EUFAR10_02-2010_199b_Santander

Ready to be archived

comment:33 Changed 8 years ago by anch

  • Summary changed from EUFAR10/02, 199b/2010, Santander to EUFAR10/02 (AIRES-CZM), 199b/2010, Santander

comment:34 Changed 8 years ago by knpa

  • Owner changed from anch to knpa
  • Status changed from accepted to assigned

Beginning archiving.

comment:35 Changed 8 years ago by knpa

Tabi data is present for this flight.

comment:36 Changed 8 years ago by knpa

Tarball created, made available to NEODC.

comment:37 Changed 8 years ago by knpa

  • Milestone set to 2010 data processing completion

Backed up by NEODC, moved to archive disk.

comment:38 Changed 8 years ago by knpa

  • Resolution set to fixed
  • Status changed from assigned to closed

Deleted workspace version.

Closing ticket.

comment:39 Changed 7 years ago by emca

  • Resolution fixed deleted
  • Status changed from closed to reopened

comment:40 Changed 7 years ago by emca

Reprocessed using APL software.
No issues found with data.

comment:41 Changed 7 years ago by emca

Reprocessed hyperspectral data ready for delivery check

comment:42 Changed 7 years ago by knpa

Delivery checking hyperspectral.

Mapped files look OK, am zipping them up as we urgently need space.

comment:43 Changed 7 years ago by knpa

Delivery check complete.

All looks fine apart from:

  • From the screenshot, e17 looks like it still has an sct error
  • The name of the igm in the example commands looks messed up - shouldn't the 30N be next to the utm part of the filename rather than the WGS84?
  • Logsheet filename incorrect (last _ should be a -)

comment:44 Changed 7 years ago by emca

Changed logsheet name. IGM files are in standard name format - maybe we should change this so it matches the mapped file names. Going to keep as they are for now for consistency.
Reprocessing full set of SCT's for eagle17 to check sct value

Last edited 7 years ago by emca (previous) (diff)

comment:45 Changed 7 years ago by emca

New sct value for eagle line 17: -0.03. Reprocessed this line and added to delivery.
Zipping this line and the remaining Hawk lines.

comment:46 Changed 7 years ago by emca

Reprocessing now complete. Data not requested by PI.

comment:47 Changed 7 years ago by adbe

Delivered reprocessed Hyperspectral to Elena Castillo-Lopez on 16/08/2012 on disk 93.

comment:48 Changed 6 years ago by knpa

Merged and deleted workspace version.

comment:49 Changed 3 years ago by dac

  • Resolution set to fixed
  • Status changed from reopened to closed

Archiving

Archived data available from NEODC (http://browse.ceda.ac.uk/browse/neodc/arsf/2010/EUFAR10_02/EUFAR10_02-2010_199b_Santander). Marking as closed.

Note: See TracTickets for help on using tickets.